# The Faiss Library

### v3

@@ 8 Conclusion
Throughout the years, Faiss continuously expanded its focus to include the most relevant vector indexing techniques from research. We plan to continue doing this to include novel quantization techniques [39], better hardware support for some indexes [63] and new indexing forms, such as associative vector memories for transformer architectures [20, 92].
@@ A.1 Code structure
Faiss is modular and includes few dependencies, so that linking it from C++ is easy. On the Python level, callback classes (ResultHanlder, InvertedLists, IDSelector) can be subclassed and wrapped in SWIG so that they can be provided to Faiss without recompiling Faiss itself.
@@ A.3 Optimization
When operations are simple enough (e.g. elementwise vector sum), the code is written in a way that the compiler can vectorize the code by itself, which often boils down to adding restrict keywords to signal that arrays are not overlapping.
@@
Selecting the $k$ closest vectors to a query vector by ranking distances on the CPU is best implemented with a min- or max-heap. On the GPU, the sequential operations involved in heap operations would similarly force the GPU into a latency-bound regime. This is the largest challenge for GPU implementation of vector search, as the time needed for the heap implementation an order of magnitude greater than all other arithmetic. To handle this, we developed an efficient GPU $k$-selection algorithm [47] that allows for ranking candidate vectors in a single pass, operating at a substantial fraction of peak possible performance per memory bandwidth limits. It relies upon heavy usage of the high-speed, large register memory on GPUs, and small-set bitonic sorting via warp shuffles with buffering techniques.
@@ A.4 Interfacing with external storage
Faiss indexes are based on simple storage classes, mainly std::vector to make copy-construction easier. The default implementation of IndexIVF is based on this storage. However, to give vector database developers more control over the storage of inverted lists, Faiss provides two lower-level APIs.
@@ A.5 How to choose an index
A.6 API index of the Faiss library
@@
  • add_with_ids($x$,$I$) add the vectors $x$, identified by the 63-bit ids $I$;
@@
  • remove_ids($I$) remove vectors with ids $I$ from the index;
  • reconstruct_batch($I$) extract the vectors with ids in $I$. It is called “reconstruct” because for most index types, the returned vectors will be approximations of the original ones.

### v4

@@ 8 Conclusion
Throughout the years, Faiss continuously expanded its focus to include the most relevant vector indexing techniques from research. We continue doing this to include novel quantization techniques [39], better hardware support for some indexes [63] and new indexing forms, such as associative vector memories for transformer architectures [20, 94].
@@ A.1 Code structure
Faiss is modular and includes few dependencies, so that linking it from C++ is easy. Callback classes (ResultHanlder, InvertedLists, IDSelector) can be subclassed to customize the indices. On the Python level, they can be wrapped in SWIG to be provided to Faiss without recompiling Faiss itself.
@@ A.3 Optimization
When operations are simple enough (e.g. elementwise vector sum), the code is written in a way that the compiler can vectorize the code by itself, which often boils down to adding restrict keywords to promise that arrays are not overlapping.
@@
Selecting the $k$ closest vectors to a query vector by ranking distances on the CPU is best implemented with a min- or max-heap. On the GPU, the sequential operations involved in heap operations would similarly force the GPU into a latency-bound regime. This is the largest challenge for GPU implementation of vector search, as the time needed for the heap implementation is an order of magnitude greater than all other arithmetic. To handle this, we developed an efficient GPU $k$-selection algorithm [47] that allows for ranking candidate vectors in a single pass, operating at a substantial fraction of peak possible performance per memory bandwidth limits. It relies upon heavy usage of the high-speed, large register memory on GPUs, and small-set bitonic sorting via warp shuffles with buffering techniques.
@@ A.4 Interfacing with external storage
Faiss indexes are based on simple storage structures, mainly std::vector to make copy-construction easier. The default implementation of IndexIVF is based on this storage. However, to give vector database developers more control over the storage of inverted lists, Faiss provides two lower-level APIs.
@@ A.5 How to choose an index
A.6 Faiss metric types
Table 3 lists the metric types supported in Faiss. They are inspired by the distance metrics in the Scipy library [88]. Only the two first metric types are supported for all index types. The others are supported only for brute-force computations and non-compressed versions of IVF and HNSW (IVFFlat and HNSWFlat).
Table 3: Faiss metric types and corresponding function in scipy.distance. Note that the Faiss names are often more terse than their scipy counterparts.
Faiss name scipy name definition comments
METRIC_INNER_PRODUCT $\langle x,y\rangle$
METRIC_L2 sqeuclidean $\|x-y\|^{2}$ The Eucliean metric can be obtained via a square root
METRIC_L1 cityblock $\sum_{i}|x_{i}-y_{i}|$
METRIC_Linf chebyshev $\max_{i}|x_{i}-y_{i}|$
METRIC_Lp minkowski $\sum_{i}|x_{i}-y_{i}|^{p}$ the parameter $p$ is set via the metric_arg parameter
METRIC_Canberra canberra $\sum_{i}|x_{i}-y_{i}|/(|x_{i}|+|y_{i}|)$
METRIC_BrayCurtis braycurtis $\sum_{i}|x_{i}-y_{i}|/\sum_{i}(|x_{i}+y_{i}|)$
METRIC_JensenShannon jensenshannon $\frac{1}{2}(\mathrm{KL}(x||m)+\mathrm{KL}(y||m))$ where $m=\frac{1}{2}(x+y)$ and $\mathrm{KL}$ is the Kullback-Leiber divergence between two distributions
METRIC_Jaccard jaccard $\sum_{i}\min(x_{i},y_{i})/\sum_{i}\max(x_{i},y_{i})$
METRIC_NaNEuclidean $\frac{|V|}{d}\sum_{i\in V}(x_{i}-y_{i})^{2}$ $V=\{i\in\{1..d\}\,\,\mathrm{st.}\,\,x_{i}\neq\mathrm{NaN}\wedge x_{i}\neq\mathrm{NaN}\}$
A.7 API index of the Faiss library
@@
  • add_with_ids($x$,II) add the vectors $x$, identified by the 63-bit ids $I$;
@@
  • remove_ids(II) remove vectors with ids $I$ from the index;
  • reconstruct_batch(II) extract the vectors with ids in $I$. It is called “reconstruct” because for most index types, the returned vectors will be approximations of the original ones.